January 10, 2011

OBSERVERS: Jill Bechtold, David Sand, Marco Pedani, Michael Rutkowski, Russell Ryan

OPERATOR: David Gonzales Huerta

Afternoon: Some high clouds. Wind around 12 mps

5:47 -- Opening the dome. We will start with Lucifer imaging.

5:51 -- Primary mirror is -3.9 C, ambient is -3.7 C. We are happy.

6:05 -- Can't open dome. Left shutter will not open. Opened manually. Bruce forgot to reprogram plc when they reassembled the left pin mechanism. Turned off breaker so it won't re-engage in the morning.

6:33 -- Dome fixed. Opening. In the background, we can not get Lucifer started.

6:41. ReadoutManager window is frozen. Rebooted several times, rebooting in engineering.

6:50. Seeing is 1.8 arcsec.

7:11. Still working on getting Lucifer up. Seeing is 2.1 arcsec.

7:57 Seeing is 0.8-1.0 arcsec. Still working on Lucifer.

Off to point and collimate at SDSS04055

John Morris went to check fibers. Checked LEDs in boxes. Went to zenith.

8:05pm Seeing .9". Went to sdss0405. Seeing 1.2"

8:11pm. Still can't restart ReadoutManager. Calling Walter in Germany; he's not answering. Calling someone else in Germany.

8:20pm Skyping with Walter.

8:32pm Rebooting Lucifer gui's.

8:38pm. John Morris is going to switch the data interface box, as suggested by Walter. Seeing 1.0" If we don't get on the air by 10pm, we're switching to LBC

9:01pm. Restarting Lucifer. Now Instrument Control Gui is not working. Telescope Control Gui not working. ReadoutManager not working.

9:06pm Restarting software. It was all off.

9:11pm We are giving up on Lucifer. Switching to LBC.

9:53 moving to SN2009ig, with the OB that Ryan altered. Running FPIA.

9:57 FPIA. Have to do manual correction since the IDL code crashed on the blue side.

10:06 estimated seeing is 1.2 arcsec.

10:19 FPIA finished. 1.3 arcsec seeing.

10:19 beginning SN2009ig in V and I. First image is 0.95

10:47 Finished OB. recollimating. 1.2 arcsec.

10:57 Correcting the co-pointing at high elevations. We are concerned about our rising target later on.

11:00 David Huertas is checking the 'bogies'. Dome made strange noise. Haunted?

11:38 One of the bogeys is making a very strange noise. We are done for the night.
Topic revision: r3 - 11 Jan 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback